# Betelgeuse, Betelgeuse, Betelgeuse, Betel-buddy? Constraints on the dynamical companion to  Orionis from HST

### v1

@@ 3.2 Our FUV Observations and configuration-specific expectations
We obtained STIS E140M spectra at four different pointings with the 0.2â0.2â aperture, which we describe below in Section 3.3. Our campaign tiled the area including and immediately surrounding Betelgeuse into quadrants, as shown in Figure 3 with the expectation that (a point source, âmas) would be in one or at most two quadrants. A detection would therefore determine which quadrant Ori B is in, thereby constraining the geometry of the orbit, independent of inferences from Betelgeuseâs inferred rotation axis.
Calculations performed in the observing proposal demonstrated that, by subtracting two HST STIS E140M spectra, one with and one without , it would be possible to spot differences at the 2% level given our proposed exposure times and anticipated S/N. Likewise, the expected km/s RV shift given the systemâs orbital configuration is within the detectable limit for our chosen grating (10âkm/s for E140M); therefore, for strong emission lines from both stars, we also anticipated an apparent doubling of lines with a separation of 0.16Ă Â in the exposure that contains (see above discussion in §2.3). It is also worth noting that Betelgeuseâs chromospheric lines are substantially broadened by turbulence (km/s, Carpenter & Robinson 1997), which is expected to exceed the broadening of the low-mass companion, but would not exceed the companionâs orbital velocity. We predicted that our campaign would be sensitive to companions , similar to the expected mass of  recovered from the high-fidelity STELLA RV (Goldberg et al., 2024) and astrometric analysis (MacLeod et al., 2025). We also predicted sensitivity to young, chromospherically active 0.6 companions, consistent with the mass suggested by the detailed long-baseline RV analysis (MacLeod et al., 2025). Moreover, may be accreting at a higher rate due to focused winds from Betelgeuse and thus may be even brighter in the FUV than its isolated, 10âMyr-old counterparts in Sco-Cen. We therefore anticipated a reasonably good chance of detecting , and if not, placing constraints on remaining mass ranges allowed by a null detection.
@@ 3.3 HST configuration
Observations for this project were obtained using the STIS instrument on HST via a Cycle 32 Directorâs Discretionary award of 4 orbits (GO/DD-17881555stsci.edu/hst-program-info/program/?program=17881; PI: M. Joyce). The observations were separated into 4, 1-orbit visits in order to ease the scheduling burden introduced by the reduced gyro mode, and also to limit the number of repeat visits required should a failure occur. The first two visits executed on 2024 November 11 with the third and fourth visits occurring on 2024 November 15 and 25, respectively.
All four visits followed the same observing plan, which utilized a 1 second imaging target acquisition with the F25ND5 attenuating filter followed by 21000 second spectroscopic exposures using the E140M grating, 0.2â0.2â aperture, and a central wavelength of 1425 Ă . The central wavelength of 1425 Ă Â was chosen to ensure that key emission features from , such as Ly- (1215.67 Ă ), C IV (1548.2 Ă ), and fluorescent H2 lines (1446 Ă , 1505 Ă ) would be captured.
Since the low-mass companion is expected to be approximately 52âmas from the center of Betelgeuse, we adopted the tiling pattern shown in Fig. 3, which was constructed using combinations of Position Target (POS TARG) offset values of 0.084â in the X and Y coordinates. The four quadrants are labeled âNâ, âEâ, âSâ, and âWâ, loosely corresponding to their ordinal position on the sky. However, since no ORIENT constraints were applied, the roll angle of Visit 04 (the âWâ quadrant in Fig. 3) was slightly different than the previous three visits. While the exact quadrant in which is expected to appear is uncertain, the tiling strategy employed here incorporates sufficient overlap to guarantee its inclusion in at least one visit, if it is present.
If the orbital plane of the companion is indeed perpendicular to the PA rotation axis (Kervella et al., 2018) consistent with the rotation axis found by Uitenbroek et al. (1998); Lobel & Dupree (2001), the companion would appear in quadrant E just outside the overlap region with quadrant S, and in the SE stack. Likewise, insofar as the rotational velocity persists and extends into the chromosphere, emission from Betelgeuse itself should be redshifted in the E quadrant by a few km/s relative to the W quadrant.
@@ 3.4 Timing
Both the LSP phase and pulsation phase are important to interpret relative velocity and flux signatures in the UV spectra. As discussed above, the timing of our observations was targeted to match the receding quadrature of , near Betelgeuseâs RV minimum on the LSP timescale, approximately halfway between the maximum and minimum brightness phases of the LSP.
@@
We compare the times of the HST observations to Betelgeuseâs light variations in Fig. 4. The two panels show 2000âd-long segments, i.e., covering about one LSP cycle with one intermediate segment between the two omitted from the plot. The 2011 observations analyzed by Carpenter et al. (2018), labeled Ayres+2010 for the observing proposal (Ayres, 2010), occurred near the LSPâs luminosity maximum on the LSP timescale, but while the pulsation was before and around minimum light. In contrast, our new observations (Joyce et al., 2024) were taken during the descending-luminosity phase of the LSP (broad behavior in the lower panel of Fig. 4, see also Fig. 5) and on the ascending-luminosity phase of the pulsation (local behavior in the lower panel of Fig. 4), just before maximum light.
@@
Fig. 5 shows the timing of these same observations relative to the variations in Betelgeuseâs RV. The 2011 observations coincided with the STELLA RV observations, although the RV data are rather sparse at that time. We can subtract the average velocity and a sine fit approximating the 2170-day LSP variation from the STELLA RV data to estimate the contribution from stellar pulsations to the photospheric velocity. We find near-zero to low positive RV excess values that increase over time during the 2011 observations, indicating that the star was near maximum expansion, and then begun to shrink (increasing the redshift of the stellar photosphere, as can be seen in the upper panel of Fig. 5). Unfortunately, there are no simultaneous RV data publicly available for the new measurements, but we can infer the pulsation phase from the lightcurve. The starâs pulsation phase during our observations corresponds to the ascending branch of the light curve (seen in the lower panel of Fig. 4). The pulsation of the star is currently dominated by the O1 mode (MacLeod et al., 2023), for which Goldberg et al. (2024) found a phase lag between the photometry and the RV close to . This, and comparison with earlier simultaneous observations, suggest that the star itself was likely shrinking and approaching maximum compression at the time of our observations (Joyce et al., 2024). Therefore, we expect Betelgeuse to be somewhat cooler at the photosphere during the 2011 observations and hotter during our measurements, due to its underlying pulsational behavior.666Interestingly, the extremely high-velocity compression following the great dimming in April 2020 did not lead to significantly hotter , possibly because of the mass ejection (see, e.g. Wasatonic 2022 who show a long baseline of , TiO strength, and , as well as discussions by Levesque & Massey 2020). However, in general the pulsation phase does track the , seen also in the Wasatonic (2022) data. Likewise, Betelgeuseâs V-band magnitude is slightly fainter during the Joyce et al. (2024) observations as compared to Ayres (2010) (âmag as compared to âmag).
@@ 4 Analysis of Observations
We now discuss the similarities and differences between spectra taken in the four quadrants in the context of the search for the putative companion. For each of the four quadrants, the two exposures per visit were stacked using the stistools777https://stistools.readthedocs.io/en/latest/ and the HASP swrapper script (Debes et al., 2024). Flux values with SNR1 were then masked out in all figures showing spectra. For consistency, we apply the same stacking and masking procedure to the Ayres (2010) observations analyzed by Carpenter et al. (2018), and the same masking procedure to the ULLYSES data which we redshift to the companionâs expected location as in Figure 2.
@@ 4.1 Comparison of the 4 quadrants
Figure 6 shows our complete observations in all four quadrants with the same colors as in Figure 3, labeled by their approximate relative position on the sky. We also show the stacked north-west (NW) and south-east (SE) âhemispheresâ, also with the same colors as Figure 3. Each line is offset for legibility. The 2011 observations presented by Carpenter et al. (2018) are shown in black. Broadly, we see no systematic difference above the noise level between any of the four quadrants. Notable emission features have approximately the same width and height in all quadrants, and are comparable to the same features in the 2011 spectrum. The 2011 spectrum is slightly higher in flux at longer wavelengths, consistent with expectations that the stellar surface is slightly cooler due to the Betelgeuseâs underlying pulsational phase, and is less noisy than our observations due to the stacking of 6 overlapping exposures compared to our 2 exposures per quadrant (4 per hemisphere).
@@ 4.2 Emission from diagnostic lines
Figure 7 highlights specific wavelength ranges containing diagnostic lines useful for inferring the presence of a companion (see Figure 2 and surrounding discussion in §2.3). These four wavelength ranges represent the most promising areas (other than Ly-, see §4.3) to detect the companion. We plot the NW (cyan) and SE (magenta) stacked hemispheres of our observations888We show halves instead of quadrants for ease in viewing, but the results are the same when considering all four quadrants, the Carpenter et al. (2018) data (black), and several observations of YSOs from ULLYSES Roman-Duval et al. (2020, 2025) at ages and masses indicated in the legend (dark navy-to-green). These YSOs have been distance (168 pc Joyce et al. 2020)999At the farther astrometric distance of 222pc (Harper et al., 2017), this would lead to a factor of reduction in flux. and extinction ( and MontargÚs et al. 2021) corrected to the parameters of the Betelgeuse system.
@@
First, we compare the shapes of emission features between the NW and SE hemispheres, and to the previous 2011 Carpenter et al. (2018) Betelgeuse observations. These three spectra represent three different system configurations: i) with emission only from Betelgeuse (likely the NW half), ii) with emission from the companion at a maximal RV shift of 45.5 km/s (likely the SE half), and iii) with emission from the companion at the same RV shift (transit) as Betelgeuse (Carpenter et al., 2018). We find no significant differences between the NW and SE halves. Further, with the Carpenter et al. (2018) data shifted to be at the approximate RV of our Betelgeuse observations, we find no differences between our data and that of Carpenter et al. (2018). All significant spectral features are the same to within noise limits. We thus report a clear non-detection of blueshifted emission features in excess of the chromospheric flux of Betelgeuse coming from a companion.
@@
Given this, we look to the ULLYSES YSO data to illustrate what upper limits we can place on the mass and properties of the companion. We rule out the presence of a M2Mâ companion (light green line), which has flux greater than that of Betelgeuse in all cases. For the 0.77Mâ and 0.91Mâ YSOs (navy lines), even their chromospheric features with the highest flux relative to that of Betelgeuse (panel d) appear too low in flux to be discerned from noise in our spectra.
@@
For the 1.1Mâ and 1.5Mâ YSOs, the presence of an RV-shifted feature in just one of our observed quadrants/halves would be difficult to disentangle from the noise (see, e.g. the CIV feature in panel d, as well as the CII doublet at 1334.5 and 1335.7 in panel a). In panels a), b), and c), for the narrow emission lines and not the disk emission features, we see that the flux of the 1.5Mâ YSO (e.g. the line near 1338.75 in panel a), or both that and the 1.1Mâ YSO (the CII lines near 1334.5-1336, H2 near 1505), meet or even slightly exceed that of all the Betelgeuse spectra.
@@
If is a 1.1-1.5Mâ YSO, and has a spectral profile similar to the example YSOs from ULLYSES, then we would rule out those masses, though the diversity in emission from YSOs renders these ULLYSES spectra imperfect templates for the spectrum of the companion. Moreover, will be embedded within the dusty circumstellar material surrounding Betelgeuse, which may obscure or otherwise reprocess emission from the companion. Rather, these individual objects are illustrative at best, and their masses estimates.
@@
While our observational campaign was designed to be maximally sensitive to a possible detection, given the diversity inherent in YSO spectra in the FUV, a non-detection is unfortunately only weakly constraining. Take, for example, the strong difference in overall flux in emission features between the 0.91Mâ and 1.1Mâ YSO spectra. Differences of up to an order of magnitude are visible (e.g. panel a around 1336.5), despite only 0.2Mâ between the mass estimates. Therefore, we cannot conclusively rule out the presence of a 1Mâ companion, though a smaller (M1Mâ) is favored.
4.3 Ly- Emission and Differences by quadrant
The Ly- (1215.67) emission line is another strong feature in the UV chromosphere of Betelgeuse as well as other stars and YSOs (Linsky, 2017; Carpenter et al., 2018; Roman-Duval et al., 2020, 2025). However, there is significant contamination of this feature from geocoronal emission from Earthâs atmosphere. The only HST UV observations of the Ly- line in a star uncontaminated by geocoronal emission are from stars with extremely high (150 km/s) radial velocities (Guinan et al., 2016; Schneider et al., 2019). The 22 km/s radial velocity of the Betelgeuse system is within the approximate width (35 km/s) of the observed Ly- line.
In an attempt to separate the source and geocoronal Ly- emission profiles, we extracted various cross-dispersion slices of the 2D spectra, including regions centered on the order traces where Ly- appears along with adjacent background regions on either side. Ly- appears in two adjacent orders in our E140M spectra, and for one of the orders the center trace spectrum shows evidence of a roughly 2.5 km s-1 shift relative to the background Ly- spectrum. However, no such signal was detected in the other order.
In the left panel of Figure 8, we plot the four quadrants of our Betelgeuse observations (same colors as Figure 3) along with the 2011 ASTRAL observations in black, shifted to the RV of Betelgeuse at the time of our 2024 observations. We see no evidence of an additional, RV-shifted Ly- line in any of the four quadrants. The Ly- emission has the same general shape and width (0.1) in all quadrants. We do observe a small shift in the center wavelengths of the emission feature â the east quadrant is offset from the west quadrant by km/s, while the north and south are offset from the west by approximately half that value. The implied orientation â with the eastern side of Betelgeuse redshifted relative to the western side â is consistent with the rotation axis inferred by Dupree et al. (1987); Gilliland & Dupree (1996); Uitenbroek et al. (1998); Lobel & Dupree (2001); Kervella et al. (2009, 2018), and others.
While large-scale convection ubiquitous in RSG envelopes (e.g. Stothers, 1972; Goldberg et al., 2022; Chiavassa et al., 2011, 2024; Ma et al., 2025) could over short timescales mimic both a changing astrometric signal (Chiavassa et al., 2020) and rotation inferred from an apparent redshift-blueshift dipole in resolved interferometric RV observations (Ma et al., 2024), convection is a stochastic process with decoherence times of order the envelope sound-crossing time (yr). The timing of our observations, on the other hand, has no correlation with the underlying convective phase or plume geometry of Betelgeuse. Therefore, the alignment of a south-westward RV dipole, persistent across many decades of observations and across many wavelengths including our 2024 observations, favors the claim that there is indeed coherent rotation of the near-surface layers.
When comparing our observations to the 2011 observations of Carpenter et al. (2018), we also find a surprising km/s offset between our median Ly- measurements and the 2011 Ly- measurement. This is the only emission feature throughout the entire spectrum that displays this offset between the 2024 and 2011 observations, including the 1302-1306 OI lines, which are also contaminated by geocoronal emission. All other features are only offset by 2.5 km/s, the RV difference in Betelgeuse between these dates, not offset by this km/s blueshift. We cannot explain this offset, but note that our 2024 data do show Ly- at approximately the same RV as other features in the spectrum.
Finally, in the right panel of Figure 8, we present our NW and SE spectra alongside several ULLYSES (Roman-Duval et al., 2020, 2025) YSOs, again distance and extinction corrected and shifted to âs RV at this phase. The YSOs plotted are different than in Figure 7; here we plot only ULLYSES YSOs observed with STIS, as COS is even more impacted by the geocoronal contamination at Ly- than STIS.
Though we were not able to extract the true shape of the Ly- emission unique to Betelgeuse, we do note that the total RV of relative to Earth is 65 km/s (the vertical orange dotted line in Figure 8), and is beyond the extent of the Ly- line. As seen in the right panel of Fig. 8, all reference YSO spectra (, , ) other than the 0.76Mâ YSO, the Ly- meet or exceed the flux of our Betelgeuse observations.
While the lack of a feature at the location of the Ly- line for the RV of is clear, we cannot firmly conclude that Ly- emission from a YSO with mass Mâ would be visible above the noise level of our 2024 spectra, as the Ly- features of the ULLYSES data are themselves dominated by geocoronal contamination. If we nonetheless expect comparable intrinsic Ly- emission from YSOs, then the lack of a second emission line here is constraining. However, without uncontaminated emission profiles, it is impossible to place an exact mass constraint using Ly-.

### v2

@@ 3.2 Our FUV Observations and configuration-specific expectations
We obtained STIS E140M spectra at four different pointings with the 0.2”0.2” aperture, which we describe below in Section 3.3. Our campaign tiled the area including and immediately surrounding Betelgeuse into quadrants, as shown in Figure 3 with the expectation that (a point source, mas) would be in one or at most two quadrants. A detection would therefore localize Ori B within a quadrant or hemisphere, thereby providing a constraint on the geometry of the orbit independent of inferences from Betelgeuse’s apparent rotation axis. In addition to the location of Betelgeuse and expected location of the companion if it is aligned with the Kervella et al. (2018) rotation axis, we indicate the diameter Point Spread Function (PSF) of STIS E140M as a horizontal bar. If the companion is present within the aperture of a given observation, then its observed image will be blurred over this range, with the majority of the flux remaining centrally concentrated. Flux from Betelgeuse itself, present in all 4 quadrants, will also be smeared. This provides additional motivation for quadrant-to-quadrant and hemisphere-to-hemisphere comparisons, rather than a finer-resolution search.
Calculations performed in the observing proposal demonstrated that, by subtracting two HST STIS E140M spectra, one with and one without , it would be possible to spot differences at the 2% level given our proposed exposure times and anticipated S/N. Likewise, the expected km/s RV shift given the system’s orbital configuration is within the detectable limit for our chosen grating (10 km/s for E140M); therefore, for strong emission lines from both stars, we also anticipated an apparent doubling of lines with a separation of 0.16Å in the exposure that contains (see above discussion in §2.3). It is also worth noting that Betelgeuse’s chromospheric lines are substantially broadened by turbulence (km/s, Carpenter & Robinson 1997), which is expected to exceed the broadening of the low-mass companion, but would not exceed the companion’s orbital velocity. We predicted that our campaign would be sensitive to companions , similar to the expected mass of recovered from the high-fidelity STELLA RV (Goldberg et al., 2024) and astrometric analysis (MacLeod et al., 2025). We also predicted sensitivity to young, chromospherically active 0.6 companions, consistent with the mass suggested by the detailed long-baseline RV analysis (MacLeod et al., 2025). Moreover, may be accreting at a higher rate due to focused winds from Betelgeuse and thus may be even brighter in the FUV than its isolated, 10 Myr-old counterparts in Sco-Cen. We therefore anticipated a reasonably good chance of detecting , and if not, placing constraints on remaining mass ranges allowed by a null detection.
@@ 3.3 HST configuration
Observations for this project were obtained using the STIS instrument on HST via a Cycle 32 Director’s Discretionary award of 4 orbits (GO/DD-17881777stsci.edu/hst-program-info/program/?program=17881; PI: M. Joyce). The observations were separated into 4, 1-orbit visits in order to ease the scheduling burden introduced by the reduced gyro mode, and also to limit the number of repeat visits required should a failure occur. The first two visits executed on 2024 November 11 with the third and fourth visits occurring on 2024 November 15 and 25, respectively.
All four visits followed the same observing plan, which utilized a 1 second imaging target acquisition with the F25ND5 attenuating filter followed by 21000 second spectroscopic exposures using using the broad-band E140M-1425 grating setting (1150-1700 Å) and 0.2”0.2” aperture. This choice ensured that key emission features from , such as Ly- (1215.67 Å), C IV (1548.2 Å), and fluorescent H2 lines (1446 Å, 1505 Å) would be captured.
Since the low-mass companion is expected to be approximately 52 mas from the center of Betelgeuse, we adopted the tiling pattern shown in Figure 3, which was constructed using combinations of Position Target (POS TARG) offset values of 0.084” in the X and Y coordinates. The four quadrants are labeled “N”, “E”, “S”, and “W”, loosely corresponding to their ordinal position on the sky. However, since no ORIENT constraints were applied, the roll angle of Visit 04 (the “W” quadrant in Figure 3) was slightly different than the previous three visits. While the exact quadrant in which is expected to appear is uncertain, the tiling strategy employed here incorporates sufficient overlap to guarantee its inclusion in at least one visit, if it is present.
If the orbital plane of the companion is indeed perpendicular to the PA rotation axis (Kervella et al., 2018) consistent with the rotation axis found by Uitenbroek et al. (1998); Lobel & Dupree (2001), the companion would appear in quadrant E just outside the overlap region with quadrant S, and in the SE stack.
@@ 3.4 Timing
Both the LSP phase and pulsation phase are important to interpret relative velocity and flux signatures in the UV spectra. As discussed above, the timing of our observations was targeted to match the receding quadrature of , near Betelgeuse’s RV minimum on the LSP timescale, approximately halfway between the maximum and minimum brightness phases of the LSP.
@@
We compare the times of the HST observations to Betelgeuse’s light variations in Figure 4. The two panels show 2000 d-long segments, i.e., covering about one LSP cycle with one intermediate segment between the two omitted from the plot. The 2011 observations analyzed by Carpenter et al. (2018), labeled Ayres+2010 for the observing proposal (Ayres, 2010), occurred near the LSP’s luminosity maximum on the LSP timescale, but while the pulsation was before and around minimum light. In contrast, our new observations (Joyce et al., 2024) were taken during the descending-luminosity phase of the LSP (broad behavior in the lower panel of Figure 4, see also Figure 5) and on the ascending-luminosity phase of the pulsation (local behavior in the lower panel of Figure 4), just before maximum light.
@@
Figure 5 shows the timing of these same observations relative to the variations in Betelgeuse’s RV. The 2011 observations coincided with the STELLA RV observations, although the RV data are rather sparse at that time. We can subtract the average velocity and a sine fit approximating the 2170-day LSP variation from the STELLA RV data to estimate the contribution from stellar pulsations to the photospheric velocity. We find near-zero to low positive RV excess values that increase over time during the 2011 observations, indicating that the star was near maximum expansion, and then begun to shrink (increasing the redshift of the stellar photosphere, as can be seen in the upper panel of Figure 5). Unfortunately, there are no simultaneous RV data publicly available for the new measurements, but we can infer the pulsation phase from the lightcurve. The star’s pulsation phase during our observations corresponds to the ascending branch of the light curve (seen in the lower panel of Figure 4). The pulsation of the star is currently dominated by the O1 mode (MacLeod et al., 2023), for which Goldberg et al. (2024) found a phase lag between the photometry and the RV close to . This, and comparison with earlier simultaneous observations, suggest that the star itself was likely shrinking and approaching maximum compression at the time of our observations (Joyce et al., 2024). Therefore, we expect Betelgeuse to be somewhat cooler at the photosphere during the 2011 observations and hotter during our measurements, due to its underlying pulsational behavior.888Interestingly, the extremely high-velocity compression following the great dimming in April 2020 did not lead to significantly hotter , possibly because of the mass ejection (see, e.g. Wasatonic 2022 who show a long baseline of , TiO strength, and , as well as discussions by Levesque & Massey 2020). However, in general the pulsation phase does track the , seen also in the Wasatonic (2022) data. Likewise, Betelgeuse’s V-band magnitude is slightly fainter during the Joyce et al. (2024) observations as compared to Ayres (2010) ( mag as compared to mag).
@@ 4 Analysis of Observations
We now discuss the similarities and differences between spectra taken in the four quadrants in the context of the search for the putative companion. For each of the four quadrants, the two exposures per visit were stacked using the stistools999https://stistools.readthedocs.io/en/latest/ and the HASP swrapper script (Debes et al., 2024). Flux values with SNR1 were then masked out in all figures showing spectra. For consistency, we apply the same stacking and masking procedure to the Ayres (2010) observations analyzed by Carpenter et al. (2018), and the same masking procedure to the ULLYSES data which we redshift to the companion’s expected location as in Figure 2.
@@ 4.1 Comparison of the 4 quadrants
Figure 6 shows our complete observations in all four quadrants with the same colors as in Figure 3, labeled by their approximate relative position on the sky. We also show the stacked north-west (NW) and south-east (SE) ‘hemispheres’, also with the same colors as Figure 3. Each line is offset for legibility. The 2011 observations presented by Carpenter et al. (2018) are shown in black. Broadly, we see no systematic difference above the noise level between any of the four quadrants. Notable emission features have approximately the same width and height in all quadrants, and are comparable to the same features in the 2011 spectrum. The 2011 spectrum is slightly higher in flux at longer wavelengths, consistent with expectations that the stellar surface is slightly cooler due to the Betelgeuse’s underlying pulsational phase, and is less noisy than our observations due to the stacking of 6 overlapping exposures compared to our 2 exposures per quadrant (4 per hemisphere).
Individual spectral features in the different quadrants do exhibit some differences from each other. Figure 7 shows the region surrounding the H2 1504.82Å line and measurement uncertainties for the four quadrants as well as the two stacked hemispheres. Differences in the line shape coming from Betelgeuse itself are to be expected, and have been observed in other spatially resolved spectroscopic observations with HST (e.g. Uitenbroek et al., 1998; Dupree et al., 2005, 2020; Granzer et al., 2022). Large-scale convection is expected to be ubiquitous in RSG envelopes (e.g. Stothers, 1972; Goldberg et al., 2022; Chiavassa et al., 2011, 2024; Ma et al., 2025), will induce large-scale asymmetries in measurements of the stellar photocenter (Chiavassa et al., 2020) and spatially resolved RV measurements (Ma et al., 2024). If the convective activity seeds chromospheric activity, this should indeed impact Betelgeuse’s emission line profiles at the quadrant-by-quadrant and hemisphere-by-hemisphere level.
Figure 7 also shows spectroscopic differences at the location of the H2 line at the companion’s expected RV, indicated by a vertical grey dotted line. While hints of a feature may appear in some quadrants, there is no clear evidence of emission lines from the companion at its expected RV offset within the measurement uncertainties in this feature. We now turn to discussing other stronger diagnostic lines.
@@ 4.2 Emission from diagnostic lines
While Ly- normally would be the most promising region to search for emission from a companion, the FUV spectrum shortward of 1235Å in Betelgeuse appears to show no flux above the noise level, potentially absorbed by Ly- damping wings in the outflowing wind. We note also that the feature near Ly- in Betelgeuse’s spectrum is primarily geocoronal (discussed in Appendix A). Given the varied nature of supergiant wind models (see van Loon 2025 for a recent review), it is difficult to assess the neutral absorption column along the line of sight to the putative companion’s orbital position. Thus we cannot determine whether the observed lack of emission in the Ly- region is due to absence of a YSO companion, or whether tangential absorption from Betelgeuse’s wind is significant enough to mask the presence of a companion with a relatively narrow Ly-alpha emission feature.
@@
Therefore, we turn to alternative wavelength regions containing promising diagnostic lines for inferring the presence of a companion (see Fig. 2 and surrounding discussion in § 2.3). Examples are shown in Fig. 8. We plot the NW (cyan) and SE (magenta) stacked hemispheres of our observations101010We show halves instead of quadrants for ease in viewing, but the results are the same when considering all four quadrants, the Carpenter et al. (2018) data (black), and several observations of YSOs from ULLYSES Roman-Duval et al. (2020, 2025) at ages and masses indicated in the legend (dark navy-to-green). These YSOs have been distance (168 pc Joyce et al. 2020)111111At the farther astrometric distance of 222pc (Harper et al., 2017), this would lead to a factor of reduction in flux. and extinction ( and , Montargès et al. 2021) corrected to the parameters of the Betelgeuse system, as detailed in §2.3.
We note that caution is needed when comparing the observed extinction applied here () to the discussion of the absorbing Hydrogen column () by O’Grady et al. (2025). Because the X-ray absorption column close to the stellar surface is unconstrained from direct X-ray observations, O’Grady et al. (2025) adopt estimates from stellar wind models of near the companion’s location. Even higher estimates of can be inferred from ALMA observations, e.g., integrating the profile determined in recent work by Dent et al. (2024). Applying the Predehl & Schmitt (1995) linear relationship between extinction and Hydrogen column for interstellar dust, , to even the low () or moderate () values of from O’Grady et al. (2025) yields implausibly high extinction values of or , respectively. In fact, the ISM relation does not apply to the partially-ionized circumstellar regime, where the dust-to-gas ratio is significantly smaller than it is in the farther-out ISM. A large hydrogen column in the star’s circumstellar wind environment could contribute to photoelectric absorption of X-rays, whereas a smaller amount of dust further out could modestly affect the FUV extinction.
@@
We compare the shapes of emission features between the NW and SE hemispheres, and to the previous 2011 Carpenter et al. (2018) Betelgeuse observations. These three spectra represent three different system configurations: i) with emission only from Betelgeuse (likely the NW half), ii) with emission from the companion at a maximal RV shift of 45.5 km/s (likely the SE half), and iii) with emission from the companion at the same RV shift (transit) as Betelgeuse (Carpenter et al., 2018). We find no significant differences between the NW and SE halves. Further, with the Carpenter et al. (2018) data shifted to be at the approximate RV of our Betelgeuse observations, we find no differences between our data and that of Carpenter et al. (2018). All significant spectral features are the same to within noise limits. We thus report a clear non-detection of blueshifted emission features in excess of the chromospheric flux of Betelgeuse coming from a companion.
@@
Given this, we look to the ULLYSES YSO data to illustrate what upper limits we can place on the mass and properties of the companion. We rule out the presence of a M2M⊙ companion (light green line), which has flux greater than that of Betelgeuse in all cases. For the 0.77M⊙ and 0.91M⊙ YSOs (navy lines), even their chromospheric features with the highest flux relative to that of Betelgeuse (panel d) appear too low in flux to be discerned from noise in our spectra.
@@
For the 1.1M⊙ and 1.5M⊙ YSOs, the presence of an RV-shifted feature in just one of our observed quadrants/halves would be difficult to disentangle from the noise (see, e.g. the CIV feature in panel d, as well as the CII doublet at 1334.5 and 1335.7 in panel a). In panels a), b), and c), for the narrow emission lines and not the disk emission features, we see that the flux of the 1.5 YSO (e.g. the line near 1338.75Å in panel a), or both that and the 1.1 YSO (the CII lines near 1334.5-1336Å, H2 near 1505Å), meet or even slightly exceed that of all the Betelgeuse spectra.
@@
If is a 1.1-1.5M⊙ YSO, and has a spectral profile similar to the example YSOs from ULLYSES, then we would rule out those masses, though the diversity in emission from YSOs renders these ULLYSES spectra imperfect templates for the spectrum of the companion. Moreover, will be embedded within the dusty circumstellar material surrounding Betelgeuse, which may obscure or otherwise reprocess emission from the companion. Rather, these individual objects are illustrative at best, and their masses estimates.
@@
While our observational campaign was designed to be maximally sensitive to a possible detection, given the diversity inherent in YSO spectra in the FUV, a non-detection is unfortunately only weakly constraining. Take, for example, the strong difference in overall flux in emission features between the 0.91M⊙ and 1.1M⊙ YSO spectra. Differences of up to an order of magnitude are visible (e.g. panel a around 1336.5Å), despite only 0.2M⊙ between the mass estimates. Therefore, we cannot conclusively rule out the presence of a 1M⊙ companion, though a smaller (M1M⊙) is favored by this analysis.